# Methods for Identifying Minimal Sufficient Statistics

### 11 Mar 2026 v1

@@ 4 Proofs
The previous proposition can be generalized to separable measurable spaces, as can be seen in Cavalcante (2025, Proposition 2.1.4).
@@
Define $S:\mathcal{X}\to\mathbb{R}^{\mathbb{N}}$ by $S(x)(n):=g_{\theta_{n}}(x)$. Then $S$ is measurable. Furthermore, it is straightforward to prove that $\sigma(S)=\vee_{n\in\mathbb{N}}\sigma(g_{\theta_{n}})$, allowing us to conclude, using Proposition 4.7.ii, that $\sigma(S)$ is minimal $\mathcal{F}$-sufficient since $\{\theta_{n}\}_{n\in\mathbb{N}}$ is dense in $(\overline{\Theta_{2}},\tau\langle\mathcal{F}\rangle)$. Next, we will prove that, for any $x,y\in\mathcal{X}$ satisfying $S(x)=S(y)$, we have $T(x)=T(y)$.
@@
Therefore, for any $x,y\in\mathcal{X}$ satisfying $S(x)=S(y)$, we have $T(x)=T(y)$, which implies, using Proposition 4.9, that there exists a measurable function $h:\mathbb{R}^{\mathbb{N}}\to\mathcal{T}$ such that $T=h(S)$.
By hypothesis, $T$ is $\mathcal{E}$-sufficient; therefore, $T$ is $\mathcal{F}$-sufficient, allowing us to conclude, using the previous equality and the fact that $S$ is minimal $\mathcal{F}$-sufficient, that $T$ is minimal $\mathcal{F}$-sufficient. Hence, $T$ is $\mathcal{E}$-sufficient and minimal $\{P_{\theta}\}_{\theta\in\overline{\Theta_{2}}}$-sufficient, allowing us to conclude, using the definition of minimal sufficiency of statistics and recalling that $\mathcal{E}\equiv\{P_{\theta}\}_{\theta\in\overline{\Theta_{2}}}$ , that $T$ is minimal $\mathcal{E}$-sufficient. ∎
@@
From the previous implications, we conclude that, for any $x,y\in\mathcal{X}$, we have $T(x)=T(y)$ iff $y\in D(x,\Theta_{0})$, which allows us to conclude, using Lemmas 4.1 and 4.2 and Method 3.1, that $T$ is minimal $\mathcal{E}$-sufficient. ∎
@@ 5 Acknowledgments
Appendix A Basic definitions
Some basic formal definitions are revisited in this section.

### 03 Apr 2026 v2

@@ 4 Proofs
The previous proposition can be generalized to separable measurable spaces, as can be seen in Cavalcante (2026, Proposition 2.1.4).
@@
The next corollary shows that the conditions required by Proposition 4.9 can be relaxed if we do not require equality at every point, which is a condition that we do not need to prove our methods. Before we prove it, we need the following lemma.
In the next lemma, $\mathfrak{B}_{\mathbb{R}}$ denotes the Borel $\sigma$-algebra of $\mathbb{R}$.
Lemma 4.1.
Let $(\mathcal{X},\Sigma,\mathbb{P})$ be a probability space with $(\mathcal{X},\Sigma)$ being an analytic Borel space, and let $(\mathcal{Y},\Sigma_{\mathcal{Y}})$ be a separable measurable space and $f:\mathcal{X}\to\mathcal{Y}$ a measurable function. If $\alpha:\mathcal{Y}\to\mathbb{R}$ is injective and $\Sigma_{\mathcal{Y}}=\alpha^{-1}(\mathfrak{B}_{\mathbb{R}})$, then there exists a measurable function $\beta:\mathbb{R}\to\mathcal{Y}$ such that $\beta\circ\alpha\circ f=f$ $\mathbb{P}$-a.e.
Proof.
Suppose that $\alpha:\mathcal{Y}\to\mathbb{R}$ is an injective function such that $\Sigma_{\mathcal{Y}}=\alpha^{-1}(\mathfrak{B}_{\mathbb{R}})$.
In the proof of Proposition 4.8, we demonstrated that our definition of analytic Borel spaces coincides with the definition of analytic spaces given by Cohn (2013). Therefore, using Corollary 8.4.3 and Lemma 8.6.1 of Cohn (2013), we can conclude that there exists $B\in\mathfrak{B}_{\mathbb{R}}$ such that $B\subseteq(\alpha\circ f)[\mathcal{X}]$ and $\mathbb{P}((\alpha\circ f)^{-1}[B])=1$.
Choose $y_{0}\in\mathcal{Y}$ and define $\beta:\mathbb{R}\to\mathcal{Y}$ by $\beta(t):=\begin{cases}\alpha^{-1}(t),&t\in B\\ y_{0},&t\notin B\end{cases}$.
Next, we will prove that $\beta$ is $(\mathfrak{B}_{\mathbb{R}},\Sigma_{\mathcal{Y}})$-measurable.
Let $E\in\Sigma_{\mathcal{Y}}$ be arbitrary. From the hypothesis, we know that there exists $A\in\mathfrak{B}_{\mathbb{R}}$ such that $E=\alpha^{-1}[A]$, allowing us to conclude that
$\beta^{-1}[E]=$ $(\beta^{-1}[E]\cap B)\cup(\beta^{-1}[E]\cap B^{c})=(\alpha[\alpha^{-1}[A]]\cap B)\cup(\beta^{-1}[E]\cap B^{c})$ (2)
$=$ $(A\cap\alpha[\mathcal{Y}]\cap B)\cup(\beta^{-1}[E]\cap B^{c})=(A\cap B)\cup(\beta^{-1}[E]\cap B^{c})$ (3)
Therefore, since $A,B\in\mathfrak{B}_{\mathbb{R}}$ and $\beta$ is constant on $B^{c}$, we infer that $\beta^{-1}[E]\in\mathfrak{B}_{\mathbb{R}}$, implying that $\beta$ is $(\mathfrak{B}_{\mathbb{R}},\Sigma_{\mathcal{Y}})$-measurable.
Since $(\alpha\circ f)(x)\in B$ for every $x\in(\alpha\circ f)^{-1}[B]$, we have $\beta(\alpha(f(x)))=f(x)$ for every $x\in(\alpha\circ f)^{-1}[B]$, implying that $\beta\circ\alpha\circ f=f$ $\mathbb{P}$-a.e., because $\mathbb{P}((\alpha\circ f)^{-1}[B])=1$.
Corollary 4.2.
Let $(\mathcal{X},\Sigma,\mathbb{P})$ be a probability space with $(\mathcal{X},\Sigma)$ being an analytic Borel space, and let $(\mathcal{Y},\Sigma_{\mathcal{Y}})$ and $(\mathcal{Z},\Sigma_{\mathcal{Z}})$ be separable measurable spaces. Suppose that $f:\mathcal{X}\to\mathcal{Y}$ and $g:\mathcal{X}\to\mathcal{Z}$ are measurable functions. If $(\forall x,y\in\mathcal{X})\big(f(x)=f(y)\Rightarrow g(x)=g(y)\big)$, then there exists a measurable function $h:\mathcal{Y}\to\mathcal{Z}$ such that $g=h\circ f$ $\mathbb{P}$-a.e.
Proof.
Suppose that $(\forall x,y\in\mathcal{X})\big(f(x)=f(y)\Rightarrow g(x)=g(y)\big)$.
By hypothesis, $(\mathcal{Z},\Sigma_{\mathcal{Z}})$ is a separable measurable space; therefore, by Corollary D.7.3.III of Cavalcante (2026), there exists an injective measurable function $\alpha:\mathcal{Z}\to\mathbb{R}$ such that $\Sigma_{\mathcal{Z}}=\alpha^{-1}(\mathfrak{B}_{\mathbb{R}})$.
The hypothesis implies that $(\forall x,y\in\mathcal{X})(f(x)=f(y)\Rightarrow(\alpha\circ g)(x)=(\alpha\circ g)(y)\big)$, allowing us to conclude, using Proposition 4.9, that there exists a measurable function $h^{\prime}:\mathcal{Y}\to\mathbb{R}$ such that
$\alpha\circ g=h^{\prime}\circ f$ (4)
By Lemma 4.1, there exists a measurable function $\beta:\mathbb{R}\to\mathcal{Z}$ such that $\beta\circ\alpha\circ g=g$ $\mathbb{P}$-a.e. Using (4), this implies that $g=h\circ f$ $\mathbb{P}$-a.e., where $h:\mathcal{Y}\to\mathcal{Z}$ given by $h:=\beta\circ h^{\prime}$ is measurable. ∎
@@
Define $S:\mathcal{X}\to\mathbb{R}^{\mathbb{N}}$ by $S(x)(n):=g_{\theta_{n}}(x)$. Then $S$ is measurable. Furthermore, it is straightforward to prove that $\sigma(S)=\vee_{n\in\mathbb{N}}\sigma(g_{\theta_{n}})$, allowing us to conclude, using Propositions 4.6 and 4.7.ii, that $S$ is minimal $\mathcal{F}$-sufficient since $\{\theta_{n}\}_{n\in\mathbb{N}}$ is dense in $(\overline{\Theta_{2}},\tau\langle\mathcal{F}\rangle)$. Next, we will prove that, for any $x,y\in\mathcal{X}$ satisfying $S(x)=S(y)$, we have $T(x)=T(y)$.
@@
Therefore, for any $x,y\in\mathcal{X}$ satisfying $S(x)=S(y)$, we have $T(x)=T(y)$, which implies, using Corollary 4.2, that there exists a measurable function $h:\mathbb{R}^{\mathbb{N}}\to\mathcal{T}$ such that $T=h(S)$ $P$-a.e. Hence, $T=h(S)$ $P_{\theta}$-a.e. for every $\theta\in\overline{\Theta_{2}}$.
By hypothesis, $T$ is $\mathcal{E}$-sufficient; therefore, $T$ is $\mathcal{F}$-sufficient, allowing us to conclude, using the previous a.e. equality and the fact that $S$ is minimal $\mathcal{F}$-sufficient, that $T$ is minimal $\mathcal{F}$-sufficient. Hence, $T$ is $\mathcal{E}$-sufficient and minimal $\{P_{\theta}\}_{\theta\in\overline{\Theta_{2}}}$-sufficient, allowing us to conclude, using the definition of minimal sufficiency of statistics and recalling that $\mathcal{E}\equiv\{P_{\theta}\}_{\theta\in\overline{\Theta_{2}}}$ , that $T$ is minimal $\mathcal{E}$-sufficient. ∎
@@
From the previous implications, we conclude that, for any $x,y\in\mathcal{X}$, we have $T(x)=T(y)$ iff $y\in D(x,\Theta_{0})$, which allows us to conclude, using Lemmas 4.2 and 4.3 and Method 3.1, that $T$ is minimal $\mathcal{E}$-sufficient. ∎
@@ 5 Acknowledgments
Appendix A Measurable-Space Background
For convenience, we collect in this appendix the measurable-space background used in Sections 3 and 4.